### v1

# Tight Bounds for Sorting Under Partial Information

Abstract
Sorting is one of the fundamental algorithmic problems in theoretical computer science. It has a natural generalization, introduced by Fredman in 1976, called sorting under partial information. The input consists of:
- a ground set $X$ of size $n$,
- a partial oracle $O_{P}$ (where partial oracle queries for any $(x_{i},x_{j})$ output whether $x_{i}\prec_{P}x_{j}$,
for some fixed partial order $P$),
- a linear oracle $O_{L}$ (where linear oracle queries for any $(x_{i},x_{j})$ output whether $x_{i}<_{L}x_{j}$,
where the linear order $L$ extends $P$)
The goal is to recover the linear order $L$ on $X$ using the fewest number of linear oracle queries.
In this problem, we measure algorithmic complexity through three metrics: the number of linear oracle queries to $O_{L}$, the number of partial oracle queries to $O_{P}$, and the time spent (the number of algorithmic instructions required to identify for which pairs $(x_{i},x_{j})$ a partial or linear oracle query is performed). Let $e(P)$ denote the number of linear extensions of $P$. Any algorithm requires worst-case $\log_{2}e(P)$ linear oracle queries to recover the linear order on $X$.
In 1984, Kahn and Saks presented the first algorithm that uses $\Theta(\log e(P))$ linear oracle queries (using $O(n^{2})$ partial oracle queries and exponential time). Since then, both the general problem and restricted variants have been consistently studied. The state-of-the-art for the general problem is by Cardinal, Fiorini, Joret, Jungers and Munro who at STOC’10 manage to separate the linear and partial oracle queries into a preprocessing and query phase. They can preprocess $P$ using $O(n^{2})$ partial oracle queries and $O(n^{2.5})$ time. Then, given $O_{L}$, they uncover the linear order on $X$ in $\Theta(\log e(P))$ linear oracle queries and $O(n+\log e(P))$ time – which is worst-case optimal in the number of linear oracle queries but not in the time spent.
We present the first algorithm that uses a subquadratic number of partial oracle queries. For any constant $c\geq 1$, our algorithm can preprocess $O_{P}$ using $O(n^{1+\frac{1}{c}})$ partial oracle queries and time. Given $O_{L}$, we uncover the linear order on $X$ using $\Theta(c\log e(P))$ linear oracle queries and time, which is worst-case optimal. We show a matching lower bound for the prepossessing also, as we show that there exist positive constants $(\alpha,\beta)$ where for any constant $c\geq 1$, any algorithm that uses at most $\alpha\cdot n^{1+\frac{1}{c}}$ partial oracle queries must use worst-case at least $\beta\cdot c\log e(P)$ linear oracle queries. Thus, we solve the problem of sorting under partial information through an algorithm that is asymptotically tight across all three metrics.
1 Introduction
Sorting is one of the fundamental algorithmic problems in theoretical computer science. It has a natural generalization, introduced in 1976 [6], called sorting under partial information. Given are a ground set $X$ of size $n$ and a partial order $P$ on $X$. The $P$ is provided as a partial oracle $O_{P}$ that for any ordered query pair $(x_{i},x_{j})$ can output in constant time whether $x_{i}\prec_{P}x_{j}$. There exists a linear oracle $O_{L}$ which specifies an unknown linear order $L$ on $X$ where $L$ extends $P$. $O_{L}$ answers queries of the form: given $x_{i}$ and $x_{j}$, is $x_{i}<_{L}x_{j}$? The goal is to recover the linear order with the fewest linear oracle queries. Let $e(P)$ denote the number of linear extensions of $P$. A query strategy is a sequence of linear oracle queries that uncovers the linear order on $P$. The log-extensions of $P$ is equal to $\log e(P)=\log_{2}e(P)$. Any query strategy, in the worst-case, needs to ask at least $\log e(P)$ linear oracle queries to recover the linear order on $P$.
Previous Work.
Fredman introduced the problem in 1976 [6, TCS’76]. He devised a query strategy that asks $\log e(P)+O(n)$ linear oracle queries. This falls short of the desired $O(\log e(P))$ bound when $\log e(P)$ is sublinear. In 1984, Kahn and Saks [8, Order’84] showed that there always exists a query which reduces the number of linear extensions by a constant fraction; yielding a query strategy that uses $O(\log e(P))$ linear oracle queries. Kahn and Kim [7, STOC’92] propose studying this problem entirely as an algorithmic problem, where in addition to counting the number linear oracle queries used, one also counts the number of algorithmic instructions required to find the next linear oracle query in the query strategy. They note that the algorithm by Kahn and Saks [8] has, in addition to its $O(\log e(P))$ linear oracle queries, exponential running time. This is because in essence their algorithm counts the linear extensions of $P$, which is a #P-complete [1] problem.
Kahn and Kim 1992 [7] propose the first polynomial time algorithm that performs $O(\log e(P))$ linear oracle queries. Their key insight is to approximate $\log e(P)$ by using the graph entropy of the incomparability graph (the undirected graph on $X$ which includes the edge $\left\{x_{i},x_{j}\right\}$ whenever they are incomparable in $P$). Their bottleneck is to compute this approximation in polynomial time via the ellipsoid algorithm. Cardinal, Fiorini, Joret, Jungers and Munro (CFJJM) [4, STOC’10] further approximate the graph entropy by computing a greedy maximum chain decomposition on the digraph induced by $P$. This way, they describe an $O(n^{2.5})$ algorithm that asks $O(\log e(P))$ linear oracle queries. Moreover, their algorithm can be separated into a preprocessing and a query phase: After spending $O(n^{2.5})$ time preprocessing $P$, the entire interaction with the linear oracle can be done in $O(\log e(P)+n)$ time using $O(\log e(P))$ linear oracle queries. See Table 1.
Preprocessing Phase Query Phase Source
Queries Time Queries Time
$O(n^{2})$ $O(n^{2})$ $\log e(P)+O(n)$ exponential [6, TCS’76]
$O(n^{2})$ $O(n^{2})$ $O(\log e(P))$ exponential [8, Order’84]
$O(n^{2})$ $O(n^{2})$ $O(\log e(P))$ $O(n^{6})$ [7, STOC’92]
$O(n^{2})$ $O(n^{2.5})$ $O(\log e(P))$ $O(\log e(P)+n)$ [4, STOC’10]
$O(n^{1+1/c})$ $O(n^{1+1/c})$ $O(c\log e(P))$ $O(c\log e(P))$ Theorem 1
$\Omega(n^{1+1/c})$ $\Omega(n^{1+1/c})$ $\frac{1}{73}c\log e(P)$ $\Omega(c\log e(P))$ Theorem 2
Table 1: Summary of previous work and our results.
Related work.
Van der Hoog, Kostityna, Löffler and Speckmann (HKLS) [12] study the problem in a restricted setting where the partial order $P$ is induced by a set of intervals (where $(x_{i},x_{j})$ are incomparable whenever their intervals $([a_{i},b_{i}],[a_{j},b_{j}])$ intersect and where otherwise $x_{i}\prec_{P}x_{j}$ whenever $b_{i}<a_{j}$). They present an algorithm that can preprocess the set of intervals in $O(n\log n)$ time such that they can compute the sorted order on $X$ in $O(\log e(P))$ linear oracle queries and algorithmic instructions. CFJRJ [3] study a variation of this problem. Let $S=(s_{1},\ldots,s_{n})$ be a set of $n$ elements and $P$ be a partial order on $S$. Let $X=(x_{1},\ldots,x_{n})$ be a set of $n$ elements and $O_{L}$ be a linear order oracle on $X$. Find a permutation $\pi$ between $S$ and $X$ such that for all $(i,j)$: $s_{i}\prec_{P}s_{j}\rightarrow\pi(s_{i})<_{L}\pi(s_{j})$, by asking questions of the form: “is $x_{i}<_{L}x_{j}$?”. They show an algorithm using $\textsc{opt}+o(\textsc{opt})+O(n)$ linear oracle queries. Roychoudhury and Yadav [10] study the problem of sorting a set into a tree. Given is a partial order $P$ on $X$ where $P$ can be induced from a tree $T_{P}$ where all edges point towards the root. Given the partial oracle $O_{P}$, the goal is to construct the tree $T_{P}$ using only queries of the form: “is $x_{i}\prec_{P}x_{j}$?” They show a randomized algorithm that produces the tree $P$ in $O(dn)$ expected time, and a deterministic algorithm to output the sorted order of $X$ in $O(\omega n+n\log n)$ time. Here $d$ is the maximal degree and $\omega$ is the width of the tree.
Contribution.
We devise the first subquadratic time algorithm that performs $O(\log e(P))$ linear oracle queries for the problem of sorting under partial information. Similar to (CFJJM) [4], we achieve separate preprocessing and query phases. For every $c\geq 1$ we show how to preprocess $P$ in $O(n^{1+\frac{1}{c}})$ time, such that given the linear oracle $O_{L}$ we can compute the sorted order on $X$ in $O(c\log e(P))$ time. We show that this result is asymptotically tight for constant $c$. Specifically, we show that there exists constants $(\alpha,\beta)$ where for any constant $c\geq 1$ any algorithm with at most $\alpha\cdot n^{1+\frac{1}{c}}$ preprocessing time and queries requires at least $\beta\cdot c\log e(P)$ linear oracle queries and time to recover the sorted order on $X$. While we use entropy-based arguments in our proofs, the algorithm itself and the lower bound construction are purely combinatorial. We prove:
Theorem 1.
For every constant $c\geq 1$, there is an algorithm that first preprocesses $P$ in $O(n^{1+1/c})$ time, and then asks $O(c\log e(P))$ linear oracle queries in $O(c\log e(P))$ time to recover the linear order on $X$, represented as a leaf-linked tree on $X$.
The constant $c$ in the theorem indicates a trade-off between the precomputation phase and the recovery phase. We complement our algorithm by an decision-tree lower bound which shows that this trade-off is a fundamental property of sorting under partial information:
Theorem 2.
Let $c$ be a constant greater or equal to $1$. There exists no deterministic algorithm that for all partial orders $P$ can preprocess $P$ in $\frac{1}{48}n^{1+\frac{1}{c}}$ partial oracle queries, such that given the oracle it can recover the linear order on $X$ in at most $\frac{1}{73}c\log e(P)$ exact and partial oracle queries.
In fact we show an even stronger lower bound, ruling out even the existence of more efficient Monte Carlo algorithms:
Theorem 3.
For $c\geq 1$ there exists a family $\mathcal{F}$ of pairs $(P,L)$ of partial orders and linear extensions such that there exists no algorithm that for more than a quarter of the $(P,L)\in\mathcal{F}$ can preprocess $P$ in $\frac{1}{48}\cdot n^{1+\frac{1}{c}}$ partial oracle queries and recover $L$ in $\frac{1}{73}\cdot c\log e(P)$ exact and partial oracle queries.
Key Ideas.
A partial order $P$ may be viewed as a directed acyclic graph (DAG) on $X$ where there exists an edge from $x_{i}$ to $x_{j}$ whenever $x_{i}\prec_{P}x_{j}$. Conversely, a DAG $D$ induces a partial order via its transitive closure $\operatorname{Tc}\left(D\right)$ which contains an edge from $x_{i}$ to $x_{j}$ whenever there is a path from $x_{i}$ to $x_{j}$ in $D$ and $x_{i}\neq x_{j}$. A chain in $P$ is any directed path in $P$ and an antichain of size $k$ is any set of $k$ pairwise incomparable elements in $P$.
(CFJJM) [4] construct a greedy chain decomposition $C_{0},C_{1},\ldots,C_{w}$ of $P$ (obtained by iteratively deleting maximum directed paths from $P$). Denote $C=\{C_{i}\}_{i=1}^{w}$ and by $E_{0}$ the set of all edges in $P$ between vertices in $C_{0}$ and $C\setminus C_{0}$. They note that this creates a partial order $P^{\prime}=\operatorname{Tc}\left(E_{0}\cup\coprod_{i=0}^{w}C_{i}\right)$ and they prove that $\log e(P^{\prime})\in O(\log e(P))$ (and thus it suffices to only consdider $P^{\prime}$ after preprocessing). In preprocessing, they spend $O(n^{2.5})$ time and $O(n^{2})$ partial oracle queries to compute the greedy chain decomposition $C_{0},C_{1},\ldots,C_{w}$. They additionally preprocess $E_{0}$ in $O(n^{2})$ time and partial oracle queries.
Given the linear oracle $O_{L}$, they pairwise merge all chains in $\{C_{i}\}_{i=1}^{w}$ into a chain $C^{*}$ in $O(\log e(P))$ time and queries through a merge-sort like procedure. The result is a partial order $P^{\prime\prime}$ where $P^{\prime\prime}=\operatorname{Tc}\left(E_{0}\cup(C_{0}\sqcup C^{*})\right)$. They merge $C_{0}$ and $C^{*}$, using the edges in $E_{0}$, in $O(n+\log e(P))$ time using an advanced merging algorithm closely tied to the entropy of bipartite graphs.
Computing a maximum chain in a DAG requires $\Omega(n^{2})$ partial oracle queries (Appendix A.3). To avoid spending at least quadratic time, we deviate from the approach of (CFJJM) by constructing a considerably sparser graph, and by computing only approximately maximal chains:
Theorem 4.
There is an algorithm that, given a partial oracle $O_{P}$ where $P$ contains an (unknown) maximum chain of length $n-k$, computes in $O(n\log n)$ queries and time a chain of length $n-2k$.
Using the above theorem as a backbone, we generalize the construction of an alternative partial order $P^{\prime}$ by computing a decomposition of $w$ chains of $P$ and $m$ antichains of $P$ where antichains have size at least $w+1$. We show that large antichains are also useful for sorting $X$ as we show:
Lemma 1.
Let a poset $(P,X)$ contain $m$ disjoint antichains of size at least $w$, then
$\log e(P)\in\Omega(m\cdot w\log w).$
In our preprocessing algorithm, we define a decomposition which we compute in $\widetilde{O}(nw^{2})$ time:
Definition 1.
For any poset $(P,X)$ and any set $V\subseteq X$ we denote by $P-V$ the graph obtained by removing all vertices corresponding to elements in $V$ from $P$. For any integer $w$, we define a partition $C_{0},A,C$ of $X$ where:
  • $C_{0}$ is an (approximately) maximum chain of $P$,
  • $A=\{A_{i}\}_{i=1}^{m}$ is a maximal set of antichains of size at least $w+1$ in $P-C_{0}$.
  • $C=\{C_{i}\}_{i=1}^{l}$ is a greedy chain decomposition of $Y=P-(C_{0}\cup A)$.
Finally, we denote by $E_{0}$ the set of edges between $C_{0}$ and $C$ in $P$ and $D^{\prime}=E_{0}\cup\coprod\limits_{i=0}^{l}C_{i}\sqcup\coprod\limits_{i=1}^{m}A_{i}$.
We prove that: (1) $D^{\prime}$ has linear size, and (2) $P^{\prime}=\operatorname{Tc}\left(D^{\prime}\right)$ satisfies $\log e(P^{\prime})\in O(\log e(P))$. To compute the decomposition of Definition 1, we use the low-width poset sorting algorithm of Daskalakis, Karp, Mossel, Riesenfeld and Verbin (DKMRV) [5] as a subroutine to prove:
Lemma 2.
There is algorithm that with $\widetilde{O}(n\,w^{2})$ time and queries partitions a poset on $n$ elements into a maximal set $A$ of antichains of size $w+1$, and a set $Y$ of at most $w$ chains.
Given our preprocessed DAG $D^{\prime}$, in the query phase, we merge all chains in $C=\{C_{i}\}_{i=1}^{l}$ into one chain $C^{*}$ with a procedure similar to merge-sort. We then merge $C_{0}$ and $C^{*}$, explicitly avoiding the advanced merging algorithm by [4]. Instead, we use a neat insertion sort strategy that is based on exponential search. Not only is this procedure considerably simpler, it requires no further preprocessing and speeds up the running time of this step from $O(\log e(P)+n)$ to $O(\log e(P))$, which is worst-case optimal even with infinite preprocessing. Finally, we show that we can insert all points in $A=\{A_{i}\}_{i=1}^{m}$ into $C_{0}\cup C^{*}$ in $O(c\cdot\log e(P))$ time using standard binary search. This last step is the bottleneck of our query phase. In Section 6 we give a matching lower bound, proving that our algorithm is asymptotically tight with respect to the number of partial order queries, linear order queries and time spent.
2 Preliminaries
A (strict) partial order $P$ on a ground set $X$ is a binary relation on $X$ that is irreflexive, asymetric and transitive. In other words, the digraph on $X$ with an edge $(x_{i},x_{j})$ for every $x_{i}\prec_{P}y_{i}$ is a transitively closed DAG. $(P,X)$ is then called a poset. When clear from context, we omit $X$ and use $P$ to denote the poset $(P,X)$. A linear order (also called total order) is a partial order with no incomparable pairs of elements. The DAG of a linear order contains a Hamilon path. A linear extension $L$ of a partial order $P$ is a linear order $<_{L}$ such that $x_{i}<_{L}x_{j}$ whenever $x_{i}\prec_{P}x_{j}$. A linear extension corresponds to a topological ordering of the DAG $P$. We assume that upper bound algorithms run on a RAM, for a pointer machine algorithm see Appendix A.2.
Input and output.
We split our algorithm into a preprocessing phase and query phase. In the preprocessing phase, our input is some partial oracle $O_{P}$ that for any ordered query pair $(i,j)$ with distinct $i$ and $j$ returns a boolean indicating whether $x_{i}\prec_{P}x_{j}$. If $x_{i}\not\prec_{P}x_{j}$ and $x_{j}\not\prec_{P}x_{i}$ then the elements are incomparable in $P$. In the query phase, our input is a linear oracle $O_{L}$ which specifies a linear order $L$ that extends $P$. For any query pair $(i,j)$ a linear oracle query outputs whether $x_{i}<_{L}x_{j}$. We measure the algorithmic complexity in the number of linear oracle and partial oracle queries and the time spent (the number of algorithmic instructions, which includes queries).
For output, we recall three previously studied problem variants: Kahn and Saks [8] and Kahn and Kim [7] output a complementing set: a set of directed edges $E\subset X\times X$ where $\forall(x_{i},x_{j})\in E$: $x_{i}<_{L}x_{j}$ and the partial order $P$ (viewed as a DAG) together with $E$ contains a Hamiltonian path. The complementing set has size $O(\log e(P))$ and can therefore be output in $O(\log e(P))$ time. Cardinal [4] have a query algorithm that runs in $O(\log e(P)+n)$ time and may thus output a linked list of $X$ in its sorted order. (HKLS) [12] study the partial order sorting problem for the restricted case where $P$ can be induced from a set of intervals (where two elements are incomparable whenever the intervals overlap). To avoid trivial $\Omega(n)$ lower bounds for reporting a linked list, they propose to output either a linked list or a balanced binary tree on $X$ (in its sorted order), by providing a pointer to either the head of the linked list or the root of the binary tree. Consider the example where $P$ is a linear order and thus $\log e(P)\in O(1)$. This definition of output allows one to preprocess $P$ in $O(n\log n)$ time and output the sorted order of $X$ in $O(1)$ time in the query phase. We preprocess any partial oracle $O_{P}$ to (given $O_{L}$) output either a complementing set, a pointer to a linked list on $X$, or the root of a balanced binary tree on $X$ in its sorted order.
Decision Trees (DT).
For our lower bound, we define decision trees that are a strictly stronger model of computation than the RAM and pointer machine model. In the preprocessing phase, a DT is a binary tree $T$ that receives as input an arbitrary partial oracle $O_{P}$. Each inner node of $T$ stores an ordered pair of integers representing the query: “$x_{i}\prec_{P}x_{j}$?”. For each inner node, its left branch corresponds to $x_{i}\not\prec_{P}x_{j}$ and its right branch to $x_{i}\prec_{P}x_{j}$. For a leaf $\ell$ in $T$, denote by $E_{\ell}$ the set of all directed edges $(x_{i},x_{j})$ where there exists a node $v$ on the root-to-leaf path to $\ell$ that stored $x_{i}\prec_{P}x_{j}$ and where the path continues right. Each leaf $\ell$ stores the DAG $Q_{\ell}=\operatorname{Tc}\left(E_{\ell}\right)$.
In the query phase, a DT is a family of trees $\{T_{Q}\}$ with a tree for each transitively closed DAG $Q$. Each tree $T_{Q}$ receives as input an arbitrary linear oracle $O_{L}$ where the linear order $L$ extends $Q$. It is a binary tree where each inner node corresponds to a query “$x_{i}<_{L}x_{j}$?”. For a leaf $s$ in the tree, denote by $E_{s}$ the edge set of all directed edges $(x_{i},x_{j})$ where there exists a node $v$ on the root-to-leaf path to $s$ that stored $x_{i}<_{L}x_{j}$ and where the path continues right. For any DT in the query phase, we require for each leaf $s$ that the graph $Q\cup E_{s}$ contains a Hamiltonian path.
Any deterministic and correct pair of preprocessing and query algorithm on a RAM or pointer machine have a corresponding pair of DTs $(T,\{T_{Q}\})$. The worst-case number of queries performed by the preprocessing algorithm is lower bounded by the height of $T$. The worst-case number of queries performed by the query phase algorithm is lower bounded by $\max_{\ell\in T}\textsc{height}(T_{Q_{\ell}})$.
Chains and antichains.
A chain in $P$ is a set $C\subseteq X$ of pairwise comparable elements. Note that $P$ defines a linear order on $C$ and that $C$ thus corresponds to a directed path in $P$. The length of a chain $C$ is $|C|$. A chain is maximum if its length is maximal among all chains. A greedy chain decomposition as introduced by (CFJJM) [4] is a partition $C$ of $X$ into chains $C_{1},\dots,C_{k}$ such that $C_{i}$ is a maximum chain on $(P,X\setminus(C_{1}\cup\dots\cup C_{i-1}))$. The complexity of computing a greedy chain decomposition is $O(n^{2.5})$ and $\Omega(n^{2})$. We may associate a partial order $P^{\prime}=\coprod_{i=1}^{k}C_{k}$ to $C$ by taking the total order induced by $P$ on each $C_{i}$; making elements of different chains incomparable. An antichain in $P$ is a set of pairwise incomparable elements. The width of $P$ is the size of the largest antichain. By Dilworth’s theorem, this equals the minimum number of chains needed to cover $P$.
Log-extensions, graph entropy, and the incomparability graph.
For any poset $P$ we denote by $e(P)$ the number of linear extensions of $P$. We define the log-extensions of $P$ to be $\log e(P)$ (henceforth, all logarithms are base $2$). To approximate the log-extensions of $P$, we define the incomparability graph $\overline{P}$ of $P$ as the undirected graph on $X$ with edges between any pair of incomparable pairs of elements. The (graph) entropy of an undirected graph $G$ was introduced by Körner 1973 [9]. For a set $S\subseteq V(G)$, let $\chi^{S}$ denote its characteristic vector, with value $1$ for every vertex in $S$ and 00 otherwise. Let:
$\operatorname{STAB}(G)=\textnormal{Convex Hull}\left\{\chi^{S}\ \Big{|}\ S\text{ is an independent set in }G\right\},$
then the entropy is defined as:
$H(G)=\min_{\lambda\in\operatorname{STAB}(G)}-\frac{1}{n}\sum_{u\in V(G)}\log(\lambda(u)).$
Note that $0\leq H(G)\leq\log n$ and $H(G^{\prime})\leq H(G)$ for any spanning subgraph $G^{\prime}\subseteq G$. The graph entropy allows us to approximate the log-extensions of $P$:
Lemma 3 (Theorem 1.1 in [7], improved to Lemma 4 in [4]).
For any partial order $P$:
$\log e(P)\leq n\cdot H(\overline{P})\leq 2\log e(P)$
Greedy chain decompositions provide yet another approximation for the log-extensions of $P$:
Theorem 5 (Theorem 2.1 in [3], rephrased as in Theorem 1 in [4]).
If $C$ is a greedy chain decomposition of $P$, then for every $\varepsilon>0$
$H(\overline{C})\leq(1+\varepsilon)H(\overline{P})+(1+\varepsilon)\log(1+1/\varepsilon),$
Setting $\varepsilon=1$ yields $H(\overline{C})\leq 2H(\overline{P})+2.$
We note that if $C=(C_{1},\ldots,C_{l})$ then a short computation shows that:
$H(\overline{C})=-\sum_{i=1}^{l}\frac{|C_{i}|}{n}\log\Big{(}\frac{|C_{i}|}{n}\Big{)}=\sum_{i=1}^{l}\frac{|C_{i}|}{n}\log\Big{(}\frac{n}{|C_{i}|}\Big{)}.$
Let $C_{0},\ldots,C_{l}$ be a greedy chain decomposition of $P$. We can store $C_{0},\ldots,C_{l}$ using only $O(n)$ space. For any partial order that is a collection of chains, there exists an asymptotically optimal merging strategy corresponding to Huffman codes:
Lemma 4 (Theorem 2 [4]).
Let $Y\subseteq X$. Let $P^{\prime}$ be the partial order induced by $P$ on $Y$. Let $C=\left\{C_{i}\right\}_{i=1}^{l}$ be a greedy chain decomposition of $P^{\prime}$. For any linear oracle $O_{L}$ extending P, consider the linear oracle $O_{L^{\prime}}$ where $L^{\prime}$ is a linear order on $Y$ induced by $L$. Given $O_{L^{\prime}}$, repeatedly merging the two smallest chains in $\left\{C_{i}\right\}_{i=1}^{l}$ (until we uncover $L^{\prime}$ on $Y$) takes $O(|Y|\cdot(1+H(\overline{P}^{\prime}))=O(\log e(P^{\prime})+|Y|)$ linear oracle queries and time. In addition, $O(\log e(P^{\prime})+|Y|)\subseteq O(\log e(P)+|Y|)$.
Finally, we use the following entropy argument by CFJJM [4]:
Lemma 5 (Lemma 5 in [4]).
Let $C$ be a maximum chain in $P$, then $n-|C|\in O(n\cdot H(\overline{P}))=O(\log e(P)).$
Finger search trees.
A finger search tree is a data structure that maintains a linearly ordered set plus an arbitrary number of fingers that each point to an element of the set. Let $n$ denote the size of the sorted set. We require the following operations:
  • (Build) Build a new tree on $n$ elements in $O(n\log n)$ time.
  • (Binary Search) Given a new element $y$, create a new finger at the largest element in the tree that is less than $y$ in $O(\log n)$ time.
  • (Exponential Search) Given a finger and a new element $y$, move the finger to largest element in tree that is less than $y$. This takes $O(1+\log d)$ time where $d$ is the number of elements between the old and new finger.
  • (Insert) Given a finger, insert a new element directly after the finger in $O(1)$ amortized time, i.e. $k$ insertions take $O(k)$ time in total, independent of any other operations performed.
The space usage is $O(n+f)$ where $f$ is the current number of fingers. Finger search trees can be implemented as level-linked red-black trees [11]. Some care has to be taken to guarantee $O(1)$ amortized insertion time after building. More complicated worst-case variants are described in [2].
3 Algorithm description
We give an overview of our algorithm. Our algorithm works in two phases. First, there is a preprocessing phase during which the algorithm interacts with the partial oracle $O_{P}$. After that, the algorithm enters the query phase and may interact with the linear oracle $O_{L}$. Our algorithm takes a parameter $c\geq 1$ specifying the trade-off between the running time of the preprocessing phase and query phase. Specifically, our preprocessing phase uses $O(n^{1+\frac{1}{c}})$ time and our query phase takes $O(c\log e(P))$ time. The respective algorithms are described in Algorithm 1 and Algorithm 2. If $O(\log e(P))$ is (super)-linear, then after the query phase we may report all points in $X$ in their sorted order in $O(\log e(P))$ time. Otherwise, we either output a complementing set of $P$, or point to the root of a leaf-linked balanced binary tree that stores the set $X$ in its sorted order. We describe how to output our tree. The complementing set may be obtained by recording all successful linear oracle queries when constructing the tree.
Algorithmic description.
During preprocessing, we compute the decomposition of Definition 1. We first remove an (approximately) maximum chain $C_{0}$ from $P$ and store it in a balanced finger search tree $T_{0}$ (in order along $C_{0}$). For some smartly chosen $w$, we compute a maximal set $A$ of antichains of size at least $w+1$ of $P-C_{0}$. This gives a set $Y=P-(C_{0}\cup A)$ which we show we obtain as a set of $w$ chains. We then partition $Y$ into a greedy chain decomposition $C=\{C_{i}^{\prime}\}_{1}^{\ell}$. Each $y\in C$ receives a finger to the maximum $x\in T_{0}$ with $x\prec_{P}y$. each $y\in C$ also receives a finger to the minimal $x\in T_{0}$ with $y\prec_{P}x$.
At query time, we merge the chains in $C$ via Lemma 4 into one chain $C^{*}$. We then merge $C^{*}$ with $C_{0}$. We do this by starting for all $y\in C^{*}$ an exponential search in $T_{0}$ from the fingers of $y$. Finally, we insert all $y\in A$ into $T_{0}$ through a regular binary search.
The main difficulty in our approach is that computing a maximum chain $C_{0}$ or a greedy chain decomposition takes $\Omega(n^{2})$ time. We avoid this lower bound in two ways: first, we compute only an approximately longest chain $C_{0}$. Second, after we have deleted large antichains from $P$, we show that we may assume that $Y=P-(C_{0}\cup A)$ is a set of at most $w$ chains. This allows us to compute a greedy chain decomposition of $Y$ in $O(nw^{2})$ time. Having merged $C$ into one long chain $C^{*}$, we need to merge $C^{*}$ and $C_{0}$ into a single chain. To avoid the quadratic preprocessing of $C_{0}$, and linear query time from the query algorithm by (CFJJM) [4], we rely upon exponential search.
Algorithm 1 Preprocessing. Input: partial oracle $O_{P}$, $c\geq 1$
1:$C_{0}\leftarrow\text{approximate maximum chain in }P$triangleright Via Theorem 4
2:$T_{0}\leftarrow\text{balanced finger search tree on }C_{0}$
3:$w\leftarrow\frac{1}{2}n^{1/(3c)}$
4:$A\leftarrow$ maximal set of antichains of size at least $w+1$ of $P-C_{0}$$\triangleright$ Via Lemma 2
5:$Y\leftarrow P-(C_{0}\cup A)$, obtained as a set of $w$ chains $\triangleright$ Via Lemma 2
6:$C\leftarrow$ greedy chain decomposition of $Y$triangleright Via  Lemma 6
7:$E_{0}\leftarrow\emptyset$
8:for all $y\in C$ do
9:     $y$ gets fingers to $\alpha_{y}=\arg\max\{x\in C_{0}\ |\ x\prec_{P}y\}$ and $\beta_{y}=\arg\min\{x\in C_{0}\ |\ y\prec_{P}x\}$
10:     $E_{0}=E_{0}\cup\{(\alpha_{y},y)$, $(y,\beta_{y})\}$
11:end for
12:$P^{\prime}=\Big{(}C_{0}\sqcup C\sqcup A\Big{)}\cup E_{0}$
13:return $(P^{\prime},T_{0},C,A)$.
Algorithm 2 Query Phase. Input: tree $T_{0}$, chain decomposition $C$, antichains $A$, oracle $O_{L}$
1:$C^{*}\leftarrow\text{chainMerge}(C)$triangleright Via Lemma 4
2:$y_{\mathrm{prev}}\leftarrow\text{nil}$
3:for all $y\in C^{*}$ in order do
4:     $\ell_{y}\leftarrow\max_{L}(y_{\mathrm{prev}},\alpha_{y})$
5:     Insert $y$ into $T_{0}$ via exponential search starting at $\ell_{y}$
6:     $y_{\mathrm{prev}}\leftarrow y$
7:end for
8:for all $y\in A$ do
9:     Insert $y$ into $T_{0}$ with binary search.
10:end for
11:return $T_{0}$
4 Analysing the Preprocessing Phase
We show that our preprocessing uses at most $O(nw^{2})$ time and queries by showing Theorem 4 and Lemmas 2 and 6 – whose application dominate the running time of Algorithm 1. In the next section, we analyse the query phase. We begin by showing Theorem 4:
See 4
Proof.
Consider an algorithm that iteratively considers two elements $y_{1},y_{2}$ that are incomparable in $P$ and removes both of them, until no such pair exists. The remaining set of vertices in $P$ is per definition a chain $C$. We claim that if the longest chain in $P$ has a maximum length of $n-k$ then $C$ has a length of at least $n-2k$. Indeed, let $C^{*}$ be a maximum chain of length $n-k$. Whenever $y_{1},y_{2}$ are incomparable, at least one of them lies outside $C^{*}$. Thus, our procedure deletes at most $k$ pairs of elements, implying $|C|\geq n-2k$.
A straight-forward implementation the above procedure takes $\Theta(n^{2})$ time and partial oracle queries. To speed this up to $O(n\log n)$, we use a sorting algorithm to find incomparable pairs. Indeed, run merge sort on $X$ using the partial oracle $P$. Whenever the merge step encounters two incomparable elements, delete both of them from $X$ and continue with the merge step as usual. When merge sort terminates, the remaining elements are linearly sorted and hence form a chain. ∎
See 2
Proof.
(DKMRV) [5] present an algorithm to partition posets of width $\leq w$ into $w$ chains. In Appendix A, we adapt the core subroutine of of their algorithm (specifically, their ‘peeling’ procedure) to instead iteratively find and delete antichains of size $w+1$. This facilitates a divide-and-conquer approach that arbitrarily splits $P$ into two parts and recursively partitions them into $w$ chains, plus some large antichains. To merge the two parts, we use our new subroutine to extract antichains of size $w+1$. After greedily extracting all large antichains, the remaining elements may be partitioned into $w$ chains by Dilworth’s theorem. ∎
Finally, given the output of Lemma 2, we show:
Lemma 6.
Given a poset $Q$ of width $\leq w$, represented via an oracle, plus a decomposition of $Q$ into $w$ chains, a greedy chain decomposition of $Q$ can be computed in $O(nw^{2})$ time and queries.
Proof.
A greedy chain decomposition can be obtained by repeatedly finding a maximum chain and deleting its vertices from the $w$ chains. Finding a single maximum chain corresponds to a longest path problem in the corresponding DAG. There, we only need to consider the edges from each element to its successor in all other chains, plus the edges on the paths corresponding to the chains. This sparsifies the DAG to contain $O(nw)$ edges and a longest path can be found in $O(nw)$ time. A similar idea was used in the ChainMerge structure of [5]. Deleting the vertices takes $O(n)$ time.
As $Q$ has width $\leq w$, the maximum chain has length $\geq n/w$, hence the number of remaining vertices decreases geometrically by a factor $(1-1/w)$ after each iteration. Thus, the total running time is $O(nw\cdot 1/(1-(1-1/w)))=O(nw^{2})$. (Note that the greedy chain decomposition might contain more than $w$ chains in general.) ∎
We can combine our observations to upper bound our preprocessing running time:
Claim 1.
The running time of Algorithm 1 is $O(n^{1+1/c})$.
Proof.
By Theorem 4, constructing $C_{0}$ takes $O(n\log n)$ time, as does building the finger search tree. By Lemma 2, we obtain $(A,Y)$ in $\widetilde{O}(nw^{2})$ time. Computing a greedy chain decomposition $C$ from $Y$ takes $O(nw^{2})$ time by Lemma 6. Finally, for each $y\in C$ we may compute the corresponding fingers $\alpha_{y}$ and $\beta_{y}$ in $O(\log n)$ time each using binary search. The for-loop therefore takes $O(n\log n)$ time and the total running time is: $\widetilde{O}(n\log n+nw^{2})=\widetilde{O}(n^{1+2c/3})=O(n^{1+1/c})$. ∎
5 Analysing the Query Phase
In this section, we show that Algorithm 2 returns a balanced binary tree on $X$ in $O(c\cdot\log e(P))$ linear oracle queries and $O(c\cdot\log e(P))$ time. Recall that our input consists of $(C_{0},A,C)$ where $C_{0}$ is an approximately maximum chain in $P$, $A$ is a maximal set of antichains of size at least $w+1$ in $P-C_{0}$ and that $C$ is a greedy chain decomposition of $Y=P-(C_{0}\cup A)$. We first note:
Claim 2.
The running time of Algorithm 2 is linear in the number of linear oracle queries, i.e. if Algorithm 2 performs $Q$ exact oracle queries, then its running time is $O(Q)$.
Proof.
The running time of Lemma 4 is linear in the number of queries performed. The same is true for the exponential search and binary search so this implies the claim. ∎
In the remainder, we upper bound the number of linear oracle queries that we perform.
Approach.
Algorithm 2 consists of three components: (1) the merging of $C$ into a chain $C^{*}$, (2) the first for-loop and (3) the second for-loop. We analyse each component separately. For (1), we immediately apply Lemma 4 which runs in $O(|Y|+\log e(P^{\prime}))\subset O(|Y|+\log e(P))$ queries. If $\log e(P)$ is $\log e(P)$ is sublinear, then $|Y|$ is the dominating term. We show that $|Y|\in O(\log e(P))$ and thus this procedure takes $O(\log e(P))$ queries. For (2), we show that merging $C^{*}$ into $C_{0}$ through exponential search uses $O(\log e(P))$ queries. Contrary to previous works that try to efficiently merge chains using the graph entropy, we use a direct counting argument. Finally for (3) we prove Lemma 1: showing that we may insert all elements in the antichains $A$ into $T_{0}$ using binary search in $O(c\log e(P))$ queries, this is the bottleneck of our approach which gives the factor $c$.
(1): Sorting $C$.
We sort all chains in $C$ into one chain $C^{*}$ using Lemma 4:
Lemma 7.
Let $(C_{0},A,C)$ be the input as specified at the section start. Given the linear oracle $O_{L}$, we can sort all chains in $C$ into a single chain $C^{*}$ in $O(\log e(P))$ linear oracle queries.
Proof.
The set $C$ is a greedy chain decomposition of $Y=P-(C_{0}\cup A)$. Let $P^{\prime}$ be the poset induced by $P$ on $Y$. We apply Lemma 4 to sort $C$ in $O(\log e(P)+|Y|)$ linear oracle queries and time. What remains is to upper bound $|Y|$. Suppose that the longest chain in $P$ has length $k$. We apply Lemma 5 to note that $k\in O(\log e(P))$. The chain $C_{0}$ has length at least $n-2k$ and thus $Y$ has at most $2k$ elements. This implies $|Y|\in O(\log e(P))$ and thus concludes the lemma. ∎
(2): The first for-loop.
Lemma 8.
Let $(C_{0},A,C)$ be the input as specified at the section start. The first for-loop of Algorithm 2 inserts all $y\in C^{*}$ into $C_{0}$ requires $O(\log e(P))$ linear oracle queries.
Proof.
For each $y\in C^{*}$, let $y$ be inserted $d_{y}$ positions after the pointer $\ell_{y}$ in $C_{0}$. I.e., if $d_{y}=1$ then $y$ is inserted immediately after $\ell_{y}$. The first for-loop takes $O(\sum\limits_{y\in C^{*}}(1+\log d_{y}))=O(|C^{*}|+\sum\limits_{y\in C^{*}}\log d_{y})=O(|Y|+\sum\limits_{y\in C^{*}}\log d_{y})$ time. In Lemma 7 we showed that $|Y|\in O(\log e(P))$, thus it remains to upper bound the second term.
Denote by $Z\subset X$ all vertices in $C_{0}\cup C^{*}$ and by $P^{\prime}$ the partial order induced by $P$ on $Z$. Denote by $L^{\prime}$ the linear order induced by $L$ on $Z$ and for all $x\in Z$ by $\pi(x)$ its index in $L^{\prime}$. To show that $\sum\limits_{y\in C^{*}}\log d_{y}\in O(\log e(P))$ we conceptually embed $Z$ onto $[1,|Z|]$ where each $x\in Z$ is a point $\pi(x)$.
For each $y\in C^{*}$, consider the interval $I_{y}:=[\pi(\ell_{y})+1,\pi(y)]$: this is an interval that contains $d_{y}$ integers and per construction of $\ell_{y}$, these intervals are pairwise disjoint. We now conceptually create a new set of points $Z^{\prime\prime}$ from $Z$. Consider simultaneously picking for every $y\in C^{*}$ an arbitrary point $z_{y}\in I_{y}$ and translating $y$ to $z_{y}$ (translating all points in $[z_{y},\pi(y)-1]$ one place to the right). Since $\ell_{y}$ is the maximum of all incoming edges in $P^{\prime}$, the result is an embedded point set $Z^{\prime\prime}$ whose left-to-right order is a linear extension $L^{\prime\prime}$ of $P^{\prime}$. Since all intervals are disjoint, the number of linear extensions $L^{\prime\prime}$ that we can create this way is $\prod\limits_{y\in C^{*}}|d_{y}|$. Thus, $\prod\limits_{y\in C^{*}}|d_{y}|\leq e(P^{\prime})\leq e(P)$ and $\sum\limits_{y\in C^{*}}\log d_{y}\in O(\log e(P))$. ∎
(3): The second for-loop.
We upper bound the running time of the second for-loop by proving Lemma 1. We show a new result involving the entropy of undirected graphs. One novelty of this approach is that it avoids dealing with the transitivity condition of partial orders.
Lemma 9.
Let $P$ be a partial order that contains a set $A$ of $m$ disjoint antichains of size at least $k$, then:
$n\cdot H(\overline{P})\geq mk\log k.$
Proof.
Let $G$ be a graph on $X$ whose edges form $m$ disjoint cliques $\left\{A_{i}\right\}_{i=1}^{m}$ of size at least $k$, corresponding to the antichains in $P$. A direct computation shows $n\cdot H(G)\geq mk\log k$. Indeed, if $\lambda\in\operatorname{STAB}(G)$, then $\sum_{x\in A_{i}}\lambda(x)\leq 1$ as an independent set in $G$ contain at most one element from $A_{i}$. By convexity of $-\log(\dots)$,
$\sum_{x\in A_{i}}-\log(\lambda(x))\geq-k\log\Big{(}\frac{\sum_{x\in A_{i}}\lambda(x)}{k}\Big{)}\geq k\log k,$
hence
$n\cdot H(G)=\min_{\lambda\in\operatorname{STAB}(G)}\sum_{x\in X}-\log(\lambda(x))\geq\min_{\lambda\in\operatorname{STAB}(G)}\sum_{i=1}^{m}\sum_{x\in A_{i}}-\log(\lambda(x))\geq\sum_{i=1}^{m}k\log k$
Finally, $H(G)\leq H(\overline{P})$ as $G$ is a spanning subgraph of $\overline{P}$. ∎
See 1
Proof.
The proof is an immediate application of Lemmas 3 and 9. ∎
Lemma 10.
The second for-loop of Algorithm 2 performs $O(c\cdot\log e(P))$ linear oracle queries.
Proof.
Inserting a single $y\in A$ into $C^{*}$ with binary search uses $O(\log n)$ queries and the for-loop of Algorithm 2 thus performs at most $O(|A|\log n)$ queries. We apply Lemma 1 with $m=|A|/(w+1)$ to note that $\log e(P)\in\Omega(|A|\log w)$. Note that $w=\frac{1}{2}n^{1+1/(3c)}$ and thus $\log_{w}(n)\in O(c)$. We now observe that the total number of queries in the second for loop of Algorithm 2 is
$O(|A|\log n)=O(\log_{w}(n)\cdot|A|\log(w))=O(c\cdot\log e(P)).\qed$
By combining Claim 1 with Lemmas 7, 8 and 10, we obtain Theorem 1.
6 Lower Bound
We fix any decision tree $T$ in the preprocessing phase of height at most $\frac{1}{48}n^{1+1/c}$ and any set of corresponding query algorithms $\{T_{Q}\}$. Our strategy is to construct a family $\{(P_{i},L_{i})\}_{i\in\mathcal{I}}$ of pairs of a partial order with a corresponding linear extension with $\log e(P_{i})=O(\frac{1}{c}n\log n)$. We will use this set to show a lower bound through the following concept:
Definition 2.
For any pair in $\{(P_{i},L_{i})\}_{i\in\mathcal{I}}$, if $O_{P_{i}}$ is the input of $T$ then the execution of $T$ ends in a leaf that stores a partial order which we denote by $Q_{i}$. We denote:
$E_{i}:=\{\textnormal{Linear orders }L_{j}\mid j\textnormal{ satisfies }Q_{j}=Q_{i}\}.$
We show that for any fixed preprocessing decision tree $T$ of height at most $\frac{1}{48}n^{1+1/c}$, for at least half of the pairs in $\{(P_{i},L_{i})\}_{i\in\mathcal{I}}$, $\log|E_{i}|\in\Omega(n\log n)=\Omega(c\cdot\log e(P_{i}))$. To show this, we show the following proposition:
Proposition 1.
For every $c\geq 3$ and $n\geq 10$, there exists a family $\left\{(P_{i},L_{i})\right\}_{i\in\mathcal{I}}$ of partial orders $P_{i}$ with a linear extension $L_{i}$ with the following properties:
  1. 1)
    $L_{i}$ is a linear extension of $P_{i}$, but not of $P_{j}$ for $j\neq i$.
  2. 2)
    $\log e(P_{i})\leq\frac{1}{c}n\log n$
  3. 3)
    If $T$ is a decision tree that performs $\frac{1}{48}n^{1+1/c}$ partial oracle queries during preprocessing, then for at least half the $i\in\mathcal{I}$, we have $Q_{i}=Q_{j}$ for $m_{i}\geq 2^{\frac{1}{72}n\log n}$ distinct $j$.
Lemma 11.
Proposition 1 implies Theorem 3.
Proof.
By Condition (1), $|E_{i}|=m_{i}$. Indeed if $Q_{j}=Q_{i}$ then per definition $L_{j}\in E_{i}$ and all $L_{j}$ are distinct. For each $i\in\mathcal{I}$, even if the set $\{(P_{i},L_{i})\}_{i\in\mathcal{I}}$ is known a priori (and thus the leaves of $T_{Q_{i}}$ contain only Hamiltonian paths corresponding to $L_{j}\in E_{i}$), the tree $T_{Q_{i}}$ contains at least $m_{i}$ leaves. Moreover, at most $22^{\lfloor\log m_{i}\rfloor-1}\leq m_{i}/$ 2 of these leaves have a depth $<\lfloor\log m_{i}\rfloor$. We may now apply Condition (3) to note that for half of the $i\in\mathcal{I}$ we have $m_{i}\geq 2^{\frac{1}{72}n\log n}$, and in particular, $\lfloor\log m_{i}\rfloor\geq\frac{1}{72}n\log n$. It follows that for at least half of the $j$ with $L_{j}\in E_{i}$, the leaf in $T_{Q_{i}}=T_{Q_{j}}$ containing $L_{j}$ has depth at least $\frac{1}{72}n\log n$. Therefore, for at least a quarter of the $j\in\mathcal{I}$, the query algorithm performs at least $\frac{1}{72}n\log n$ queries before outputting $L_{j}$. By Condition (2), $\frac{1}{72}n\log n\geq\frac{1}{72}\cdot c\log e(P_{i})$. ∎
The Family
Given $c$ and $n$, let $w=n^{1/c}$ and $h=n/(2w)$. For simplicity, assume that $w$ and $h$ are both integers. We first give our intuitive definition (see also Figure 1). We call the first $n/2$ elements the Essentials and use them to build $w$ chains of length $h$. Each $P_{i}$ can then be generated by arbitrarily placing the remaining $n/2$ elements (Leftovers) between any two Essentials. All Leftovers $x_{\ell}$ between any two consecutive Essentials $(x_{j+kw},x_{j+(k+1)w})$ are linearly ordered by $\ell$. Formally, we define $\{(P_{i},L_{i})\}$ to consist of all partial orders $P_{i}$ with the following properties
  1. a)
    $P_{i}$ is the disjoint union of $w$ chains.
  2. b)
    (Essentials) for each $j\in[w]$, the elements $\{x_{j},x_{j+w},\dots,x_{j+(h-1)w}\}$ appear on the $j$-th chain in that order. $x_{j}$ is the minimum and $x_{j+(h-1)w}$ is the maximum element of the $j$-th chain (this creates $w$ chains where each chain has $h$ elements from the first $n/2$ elements of $X$).
  3. c)
    (Leftovers) for each $\ell>n/2$ there exists a unique pair $(j,k)\in[w]\times[0,h-2]$ with $x_{j+kw}\prec_{P_{i}}x_{\ell}\prec_{P_{i}}x_{j+(k+1)w}$ All $x_{\ell}$ with $x_{j+kw}\prec_{P_{i}}x_{\ell}\prec_{P_{i}}x_{j+(k+1)w}$ are linearly ordered by $\ell$.
  4. d)
    We define $L_{i}$ as the concatenation of the $w$ chains (connecting $x_{j+(h-1)w}$ to $x_{j+1}$).
Figure 1: Our family of pairs of partial orders with linear extensions $\{(P_{i},L_{i})\}$ can be constructed in two stages. (a): Step 1: partition the vertices $x_{i}$ for $i\leq n/2$ into $w$ equal-size chains. (b) For all $\ell>n/2$, arbitrarily assign $\ell$ to lie between any two connected vertices of the previous construction. All $x_{\ell}$ that get assigned the same pair $(x_{j+kw},x_{j+(k+1)w})$ are linearly ordered by $\ell$. The corresponding linear order $L_{i}$ is obtained by adding the red edges.
Lemma 12.
The family $\{(P_{i},L_{i})\}_{i\in\mathcal{I}}$ has Properties 1) and 2) of Proposition 1.
Proof.
First we show Property 1). Given the linear extension $L_{i}$ of a partial order $P_{i}$, we can uniquely recover $P_{i}$ as follows: split the chain $L$ into $w$ chains at $x_{1},x_{2},\dots,x_{w}$ (making these the head of each chain). This uniquely recovers the $w$ chains that are $P_{i}$ and thus for all $i,j$: $L_{i}\neq L_{j}$.
To show Property 2), we note that by construction any $P_{i}$ is the disjoint union of $w$ chains. We use the definition of graph entropy, setting $\lambda\equiv 1/w$ to note that $H(\overline{P_{i}})\leq\log w=\frac{1}{c}\log n$. By Lemma 3, $\log e(P_{i})\leq nH(\overline{P_{i}})=\frac{1}{c}n\log n$. ∎
The third property
What remains is to show that $\{(P_{i},L_{i})\}_{i\in\mathcal{I}}$ satisfies Property 3). Recall that $T$ is our decision tree corresponding to the preprocessing phase. Suppose that it performed at most $\frac{1}{48}n^{1+1/c}$ partial oracle queries during preprocessing. We say a partial oracle query “is $x_{j}\prec_{P}x_{i}$?” succeeds if the oracle returns “yes”.
Definition 3.
For any decision tree $T$ of the preprocessing phase and partial order $P_{i}$, we define three conditions after running $T$ with $O_{P_{i}}$ as input:
  • An elements $x_{j}$ is interesting if $j>n/2$,
  • an interesting element is weak if at most $w/4$ queries are performed on it, and
  • an interesting element is hidden if no queries involving it succeed.
Consider the probability space defined by taking a uniformly random $P_{i}$ and running $T$ with $O_{P_{i}}$ as input. We show that our set of partial orders has Property 3) through showing four claims:
Claim 3.
For any $P_{i}$ where $O_{P_{i}}$ is the input of $T$, there are always at least $n/3$ weak elements.
Proof.
Each of the at most $\frac{1}{48}n^{1+1/c}$ queries involve at most two interesting elements. Thus, there are at most $\frac{1}{24}n^{1+1/c}/(w/4)=\frac{1}{6}n$ non-weak interesting elements. The remaining $n/2-n/6$ interesting elements are weak. ∎
Claim 4.
Let $P_{i}$ be sampled uniformly at random. $\forall x_{j}$ with $j>\frac{n}{2}$: $\mathrm{Pr}\left[x_{j}\text{ is weak and not hidden}\right]\leq\frac{1}{4}.$
Proof.
If we fix the partial order of all elements except $x_{j}$, then $x_{j}$ is equally likely to be in any of the $w$ chains, as by Condition (d) in the definition of $\left\{(P_{i},L_{i})\right\}_{i\in\mathcal{I}}$, there are exactly $h-1$ options to place $x_{j}$ on a particular chain (in between each pair of Essentials on the chain). Each query involving $x_{j}$, and $x_{k}$, succeeds only if $x_{j}$ is on the same chain as $x_{k}$. Thus
$\mathrm{Pr}\left[x_{j}\text{ is not hidden after the first $w/4$ queries involving $x_{j}$}\right]\leq\frac{w/4}{w}=\frac{1}{4}.$
Intuitively, this corresponds to trying to find a uniformly random ball among $w$ bins by looking at $\leq w/4$ bins. The best you can do is look at $w/4$ distinct bins, then you find the ball with probability $1/4$. The claim now follows as [$x_{j}$ is weak and not hidden] implies [$x_{j}$ is not hidden after the first $w/4$ queries involving $x_{j}$]. ∎
Claim 5.
Let $P_{i}$ be sampled uniformly at random. Let $H$ denote the number of hidden elements after running $T$ with $O_{P_{i}}$ as input. Then $\mathrm{Pr}\left[H\geq n/12\right]\geq 1/2$.
Proof.
4 shows that the expected number of weak non-hidden elements is $\leq n/8$. By Markov’s inequality,
$\mathrm{Pr}\left[\text{\#weak non-hidden elements}\geq n/4\right]\leq\frac{1}{2}.$
By 3, there are $\geq n/3$ weak elements. If at most $n/4$ of them are non-hidden, then at least $n/12$ of them are hidden. ∎
Claim 6.
Suppose that running $T$ with $O_{P_{i}}$ as input ends with $H$ hidden elements and at a leaf that stores $Q_{i}$. Then, there exist at least $(h-1)^{H}$ distinct $P_{j}$ compatible with the partial order $Q_{i}$.
Proof.
[…]
Concluding our argument.
We now show that $\left\{(P_{i},L_{i})\right\}_{i\in\mathcal{I}}$ satisfies Condition (3) of Proposition 1. By 5, for at least half the $P_{i}$, we have $H\geq n/24$. For those $P_{i}$, by 6, the number of $P_{j}$ compatible with the query results is at least $(h-1)^{n/24}=2^{n\log(h-1)/24}$. Finally, $c\geq 3$ and $n\geq 10$ imply $\log(h-1)\geq\log(n)/3$. Thus, by Lemma 11, we may conclude:
See 3
Appendix A Appendix
A.1 Decomposing a poset into chains and antichains
Recall that a poset has width $\leq w$ if and only if it can be decomposed into $w$ chains. By Dilworth’s theorem, this is equivalent to there being no antichain of size $w+1$. We now explain how to modify the merge sort-like algorithm for sorting posets of width $\leq w$ of (DKMRV) [5] to work in general posets by finding and deleting antichains of size $w+1$. The core of the (DKMRV) algorithm is the Peeling procedure, which consists of $w$ peeling iterations:
Theorem 6 (PeelingIteration, Theorem 3.5 in [5]).
Let $P$ be a poset of width $\leq w$ on a set $X$ of size $n$. Given an oracle $O_{P}$ and a decomposition of $P$ into $w+1$ chains, there is an algorithm that returns a decomposition of $P$ into $w$ chains in $O(nw)$ time.
To apply this peeling approach to posets of arbitrary width, we introduce a new procedure that efficiently finds and deletes large antichains.
Lemma 13 (AntichainExtraction).
Let $P$ be a poset on a set $X$ of size $n$. Given an oracle $O_{P}$ and a decomposition of $P$ into $w+1$ chains, there is an algorithm that returns a decomposition of $P$ into $w+1$ chains and a maximal set $A$ of antichains of size $w+1$, in $O(nw)$ time. (Note that the $w+1$ chains then form a decomposition of a poset of width $\leq w$ by maximality of $A$.)
We defer the description of AntichainExtraction to later. By alternating antichain extraction and peeling iterations $w$ times, we can turn a chain decomposition of $P$ with $2w$ chains into one with $w$ chains and zero or more antichains of size $w+1$, see Algorithm 3. This can be thought of as a generalization of the Peeling procedure of (DKMRV) [5]. By using this as a subroutine, we can then mimic merge sort, see Algorithm 4. We can now prove:
See 2
Proof.
We first analyze the running time of Algorithm 3. A single antichain extraction or peeling iteration takes $O(nw)$ time. As we perform $w$ of these, this takes $O(nw^{2})$ time in total.
Now we analyze Algorithm 4. The merge sort recursion has depth $\log(n/w)$, and each recursion level takes $O(nw^{2})$ time, hence the total running time is $O(nw^{2}\log(n/w))$. ∎
Algorithm 3 AntichainPeeling
Input: partial oracle $O_{P}$, chain decomposition $\left\{C_{i}\right\}_{i=1}^{2w}$ of $P$
Output: a decomposition of $P$ into $w$ chains and zero or more antichains of size $w+1$
1:$A\leftarrow\left\{\right\}$
2:for $i=w,\dots,1$ do
3:     $(C_{i}^{\prime},\dots,C_{i+w}^{\prime}),A^{\prime}\leftarrow\textsc{AntichainExtraction}(O_{P},(C_{i},\dots,C_{i+w}))$triangleright Lemma 13
4:     $C_{i}^{\prime\prime},\dots,C_{i+w-1}^{\prime\prime}\leftarrow\textsc{PeelingIteration}(O_{P},(C_{i}^{\prime},\dots,C_{i+w}^{\prime}))$triangleright Theorem 6
5:     $A\leftarrow A\cup A^{\prime}$
6:     $C_{i},\dots,C_{i+w-1}\leftarrow C_{i}^{\prime\prime},\dots,C_{i+w-1}^{\prime\prime}$
7:     $C_{i+w}\leftarrow\left\{\right\}$
8:end for
9:return $(C_{1},\dots,C_{w}),A$
Algorithm 4 Chain-Antichain Merge Sort.
Input: partial oracle $O_{P}$, set $X$
Output: a decomposition of $P$ on $X$ into $w$ chains and zero or more antichains of size $w+1$
1:procedure ChainAntichainMergeSort($O_{P},X$)
2:     if $|X|\leq w$ then
3:         $C\leftarrow$ elements of $X$ as size-1 chains.
4:         return $C,\left\{\right\}$triangleright No antichains
5:     end if
6:     $X_{1}\sqcup X_{2}\leftarrow X$triangleright Split $X$ into equally sized parts
7:     $C^{\prime},A^{\prime}\leftarrow$ ChainAntichainMergeSort($O_{P},X_{1}$)
8:     $C^{\prime\prime},A^{\prime\prime}\leftarrow$ ChainAntichainMergeSort($O_{P},X_{2}$)
9:     $\triangleright$ Now $(C^{\prime}\cup C^{\prime\prime})$ is a decomposition of $X-A^{\prime}-A^{\prime\prime}$ into $2w$ chains.
10:     $C^{\prime\prime\prime},A^{\prime\prime\prime}\leftarrow$ AntichainPeeling$(O_{P},C^{\prime}\cup C^{\prime\prime})$triangleright Algorithm 3
11:     return $C^{\prime\prime\prime},A^{\prime}\cup A^{\prime\prime}\cup A^{\prime\prime\prime}$
12:end procedure
Antichain Extraction.
Let us now show Lemma 13. Since $P$ is decomposed into $w+1$ chains, any antichain of size $w+1$ needs to contain exactly one element from each such chain. If the maximum elements of all $w+1$ chains are pairwise incomparable, then they form an antichain. Otherwise, there are two comparable maximum elements $x,y$ of different chains. Without loss of generality, let $x\prec_{P}y$, then $x^{\prime}\prec_{P}y^{\prime}$ for any element $x^{\prime}$ in the same chain as $x$, so $y$ cannot lie in an antichain of size $w+1$. The resulting algorithm is described in Algorithm 5. A naive implementation which compares all $O(w^{2})$ pairs of maximum elements at each step takes $O(nw^{2})$ time. This gets reduced to $O(nw)$ by comparing each such pair only once, which proves Lemma 13.
Algorithm 5 Antichain Extraction (heavily inspired by Figure 3 of [5])
Input: partial oracle $O_{P}$, chain decomposition $\left\{C_{i}\right\}_{i=1}^{w+1}$ of $P$
Output: a decomposition of $P$ into $w+1$ chains and a maximal set of antichain of size $w+1$
1:for $i=1,\dots,w+1$ do
2:     $C_{i}^{\prime}\leftarrow$ copy of $C_{i}$
3:end for
4:$A\leftarrow\left\{\right\}$
5:while all $C_{i}^{\prime}$ are non-empty. do
6:     if there is a pair $(x,y)$ of maximum elements $x\in C_{i}^{\prime}$, $y\in C_{j}^{\prime}$, such that $x\prec_{P}y$ then
7:         delete $y$ from $C_{j}^{\prime}$
8:     else
9:         delete the maximum element from each $C_{i}^{\prime}$
10:         add those $w+1$ elements as an antichain to $A$
11:     end if
12:end while
13:delete the elements in $A$ from their respective $C_{i}$
14:return $(C_{1},\dots,C_{w+1}),A$
A.2 RAM model vs Pointer Machine
Our algorithms as presented run in the RAM model of computation. Here, we explain how to match the bounds of Theorem 1 on a pointer machine.
For the preprocessing phase, we use a standard simulation technique: Any RAM model algorithm with running time $O(T)$ may be simulated on a pointer machine in $O(T\log T)$ time by replacing arrays with balanced search trees. Our analysis shows that the preprocessing phase runs in $\widetilde{O}(n^{1+2c/3})$ time in the RAM model, hence the simulation on a pointer machine takes $\widetilde{O}(n^{1+2c/3}\cdot\log n)\subseteq O(n^{1+1/c})$ time.
For the query phase, this simulation technique would only yield $O(\log e(P)\log\log e(P))$ running time. To avoid the additional log factor, we instead note that our query phase runs on a pointer machine as is. Indeed, the merge sort step in Algorithm 2 can be implemented with linked lists, and finger search trees work as-is on a pointer machine. Hence, our query phase runs in $O(\log e(P))$ on a pointer machine too.
A.3 A Quadratic Lower Bound for Finding Maximum Chains
In the introduction we state that computing an (exact) maximum chain requires $\Omega(n^{2})$ partial oracle queries. Let us now prove this. As the first chain in a greedy chain decomposition is a maximum chain, such a lower bound automatically applies to computing a greedy chain decomposition too.
If $P$ consists of two comparable elements, $x\prec_{P}y$, and all other pairs are incomparable, then clearly any algorithm needs to spend $\Omega(n^{2})$ partial oracle queries to find the unique maximum chain $(x,y)$. However, one might argue that this example is a bit artificial from a sorting perspective. Indeed, as $\log(e(P))=\Theta(n\log n)$, any standard sorting algorithm can sort $P$ in $O(\log e(P))$ linear oracle queries and time, so there is no need to compute a maximum chain. To avoid this trivial case, we describe a more general family of partial orders.
Lemma 14.
For every $20<k\leq n-$ 2, there is a family of partial orders $P$ with a unique maximum chain of length $n-k$, such that any algorithm needs to spend $\Omega(k^{2})$ time to find the maximum chain.
Proof.
We define $P$ as follows: Let $A$ consists of $k+2$ elements and $B$ consist of $n-k-2$ elements. Let $x,y\in A$ be distinct. Put $x\prec_{P}y$ and make all other pairs of elements in $A$ incomparable. Let the elements in $B$ form a chain. Put $a\prec_{P}b$ for every $a\in A,b\in B$.
$P$ has a unique maximum chain $\left\{x,y\right\}\cup B$, of length $2+|B|=n-k$. Any algorithm that finds this chain needs to find $(x,y)$ among all $\Theta(k^{2})$ pairs of elements in $A$. This requires $\Omega(k^{2})$ partial oracle queries. ∎
If the maximum chain should be output in sorted order, then an additional $\Omega(n\log n)$ lower bound for sorting applies, hence the total lower bound is $\Omega(n\log n+k^{2})$. Using our techniques, we can match this bound algorithmically.
Lemma 15.
Let $P$ be a partial order with a maximum chain of length $n-k$. Then, a maximum chain in $P$ can be computed in $O(n\log n+k^{2})$ time and partial oracle queries.
Proof sketch.
With Theorem 4, we can find a chain $C$ of length $n-2k$ in $O(n\log n)$ time. For the remaining $2k$ elements, we find their predecessor and successor in $C$ in $O(k\log n)$ time. To find a maximum chain, it suffices to consider (1) these edges from a predecessor / to a successor in $C$, (2) the edges in the path formed by $C$, and (3) the edges between any pair of elements not in $C$. This sparsifies $P$ into a DAG with $n$ vertices and $O(n+k^{2})$ edges. There, a longest path can be found in $O(n+k^{2})$ time. ∎